# Computation of attractor dimension and maximal sums of Lyapunov exponents using polynomial optimization

### v1

@@ 3 Computational examples
We now present computational examples in which our methods based on SOS optimization are applied to particular ODEs. It has been conjectured that, for all systems with equilibria embedded in chaotic attractors, the supremum in the formula for the Lyapunov dimension 18 is attained on an equilibrium [9, 22]. This indeed is the case for many simple systems, including Lorenz’s 1963 model [23]. In order to demonstrate the power of our methods, we focus on two example systems for which we know there are no equilibria.
@@
For all of our examples, the code to replicate the results, implemented in Julia, is available at https://github.com/jeremypparker/Parker_Goluskin_LEs. We used SumOfSquares.jl [20] to reformulate SOS optimization problems as SDPs, along with SymbolicWedderburn.jl [15] for symmetry exploitation. To solve the resulting SDPs, we primarily used Mosek [1]. Certain SDPs were also solved using Clarabel.jl [14] and Hypatia.jl [5], which gave very similar results with longer runtimes. In applications of the sphere projection approach, the bounds $B$ can be minimized as the objective of the SDPs, and the bounds we report are the numerical optima returned by Mosek. In applications of the shifted spectrum approach, the value of $B$ must be fixed during each SDP solution because it multiplies coefficients of $V$ in the SOS constraints, and the coefficients to be optimized over can appear only linearly. In such computations, if there exists a polynomial $V$ satisfying the SOS constraints, then the SDP will have a nonempty constraint set, and the SDP solver should report feasibility. We thus perform a bisection search in $B$ to find the boundary between large $B$ values where the SDPs are feasible and small $B$ values where they are infeasible. Close to this boundary the numerical conditioning of the SDPs becomes increasingly poor, making it difficult to precisely identify the infimum over admissible $B$. Confirming that near-optimal bounds are valid calls for rigorous and/or multiple-precision computations [12, 37], but for simplicity we have used only double precision arithmetic in our examples. We call a bound ‘feasible’ if Mosek reports a dual objective value of less than $10^{-6}$, ‘infeasible’ if the final dual objective value is greater than $1$ and ‘indeterminate’ otherwise, but we stress that these computations are not rigorously validated.
@@
To confirm that accuracy of our computed bounds, we computed various period orbits and the LE sums along each of these orbits. We used the Tsit5 solver from the OrdinaryDiffEq.jl package [42] to solve integrate the ODE systems, find periodic orbits and compute LEs. For a given periodic orbit of period $T$, the LEs $\mu_{k}$ were found using the eigenvalues of the Jacobian $D\phi^{T}$. To compute the sums $M_{k}$, we simply sum the $\mu_{k}$ in descending order.
@@ 3.1 The Duffing Oscillator
The Duffing oscillator is a second-order nonautonomous ODE governing amplitude of oscillations that are damped and sinusoidally forced. Different versions exist, but we choose
$\ddot{A}+\delta\dot{A}+\beta A+\alpha A^{3}=\gamma\cos{\omega t}$ (107)
with parameters $(\alpha,\beta,\gamma,\delta,\omega)=(1,-1,0.3,0.2,1)$. We numerically estimate the Lyapunov exponents on the chaotic attractor to be approximately $0.158$ and 0.358-0.358, which gives a local Lyapunov dimension of $d_{L}\approx 1.441$. Kuznetsov and Reitmann [18] derive a rigorous bound on the Lyapunov dimension of the global attractor in the Duffing oscillator which, for these parameter values, gives a bound $d_{L}\leq 1.9910$. Note that we have not defined LEs or Lyapunov dimension for nonautonomous systems. This is straightforward to do, but instead we convert 107 to an autonomous system.
The ODE 107 is not a polynomial system, as required for the formulations presented in section 2, but it can be transformed into a polynomial system by a standard trick (cf. Parker et al. [38]). We let $x_{1}=A$ and $x_{2}=\dot{A}$, and we introduce $x_{3}=\sin{\omega t}$ and $x_{4}=\cos{\omega t}$. Then the $x(t)$ vector evolves according to
$\frac{\mathrm{d}x}{\mathrm{d}t}=\begin{bmatrix}x_{2}\\ -\delta x_{2}-\beta x_{1}-\alpha x_{1}^{3}+\gamma x_{4}\\ \omega x_{4}\\ -\omega x_{3}\end{bmatrix}$. (108)
Recalling that $x_{3}$ and $x_{4}$ are sine and cosine of the same argument, we are interested only in dynamics on the manifold
$\mathcal{B}=\{x\in\mathbb{R}^{4}\,|\,x_{3}^{2}+x_{4}^{2}=1\}.$ (109)
Since only odd powers of $x$ appear in 108, the right-hand side is equivariant under the simple sign symmetry $\Lambda=-I$, and $\mathcal{B}$ is invariant, so we can also impose that our polynomials $V$ are invariant under the same symmetry. This is exploited in our code using the methods described in section 2.5.
For the system 108, the Jacobian is
@@
From this, we compute the additive compound matrices
@@
$\text{and}\quad Df^{[4]}(x)$ $=-\delta.$ (113)
In this case, the fourth additive compound matrix—the trace of the Jacobian—is independent of $x$, and therefore either of our methods gives the sharp result that $M_{4}=-\delta$ on any orbit. This implies that the dimension $d_{L}$ is optimized on the same orbit as the optimizer for $M_{3}$ and, given a sharp bound for $M_{3}$, proposition 2.7 will give a sharp bound for $d_{L}$. Furthermore, since there are no equilibria, every trajectory must have a zero LE associated with the flow, and an additional vanishing LE corresponding to amplitude change in the $x_{3}$ and $x_{4}$ variables. Therefore $M_{1}=M_{2}=M_{3}$ for all orbits of interest, and it suffices to compute bounds on $M_{1}$ only in order to deduce bounds for $M_{2}$, $M_{3}$ and $d_{L}$. Nevertheless, to test our methods we computed bounds using our formulations for $M_{1}$, $M_{2}$, $M_{3}$ and $d_{L}$.
To formulate SOS optimization problems, we must choose finite dimensional spaces for the polynomials to be optimized over. In all cases we choose a maximum degree of $V$ and then construct a basis from symmetry-invariant monomials. For each SOS constraint, the restriction that $h_{j}(x)=0$, where $h_{j}=x_{3}^{2}+x_{4}^{2}-1$, is enforced via additional unknown polynomials $\rho_{j}$, whose maximum degrees are chosen so that $h_{j}\rho_{j}$ matches the maximum degree of other terms in the expression. See the accompanying code for full details.
For the sphere projection methods proposition 2.15 and proposition 2.16, computational results are reported in tables 1 and 2 respectively. In both cases, clearly converged bounds are achieved with $V$ of degree 4 in $x$ and 2 in $z^{k}$. For the shifted spectrum methods proposition 2.17 and proposition 2.18, numerical errors are large when the degree of $V$ in $x$ is 4 or smaller, so we let $V$ have degree $6$ in $x$. In this case, we find feasible bounds $M_{1}\leq 0.887898$ and $d_{L}\leq 3.8162$, and we find infeasible bounds when each final digit is decreased by 1. These bounds perfectly match those found via the sphere projection methods.
$d_{x}$ $d_{z}$ $k=1$ $k=2$ $k=3$ $k=4$
2 0 27.3625 31.8231 29.7823 -0.2000
2 2 14.8295 15.5894 13.9366 -0.2000
4 0 1.3666 1.3668 1.3666 -0.2000
4 2 0.8879 0.8880 0.8879 -0.2000
4 4 0.8879 0.8879 0.8879 -0.2000
@@
Table 1: Upper bounds on the sum $M_{k}$ of the $k$ leading LEs, computed by solving the SOS optimization problem in proposition 2.15 for the Duffing oscillator in the variables of 108. Polynomial degrees are at most $d_{x}$ in $x$ and $d_{z}$ in $z^{k}$. Tabulated values are rounded up to the precision shown, and the underlines indicate the digits that agree (after rounding) with the values computed by numerical integration for the periodic orbit of figure 1. On this orbit $M_{4}=0.2$ and $M_{1}=M_{2}=M_{3}\approx 0.88785$.
$d_{x}$ $d_{z}$ Bound
0 0 4.0000
0 2 4.0000
0 4 4.0000
2 0 3.9912
2 2 3.9774
2 4 3.9614
4 0 3.8724
4 2 3.8162
4 4 3.8162
6 0 3.8555
6 2 3.8162
Table 2: Upper bounds on the Lyapunov dimension $d_{L}$, computed by solving the SOS optimization problem in proposition 2.16 with $k=3$ for the Duffing oscillator in the variables of 108. Polynomial degrees are at most $d_{x}$ in $x$ and $d_{z}$ in $z^{k}$ and $z^{k+1}$. Tabulated values are rounded up to the precision shown, and underlines indicate digits that agree (after rounding) with the value $d_{L}\approx 3.81615$ computed by numerical integration for the periodic orbit of figure 1.
Figure 1: A chaotic trajectory (grey) in the system 108, and the unstable periodic orbit of period $2\pi$ discussed in the text (blue).
To find an orbit that maximises the leading LE $M_{1}$, and consequently $d_{L}$ also, we follow a strategy similar to that of Tobasco et al. [46]. We restrict to a Poincaré section of $\mathcal{B}$ with $x_{3}=0$ and $x_{4}=1$, then we attempt to find the global minimum of the expression
$B-\Phi_{k}-f\cdot D_{x}V-\ell_{k}\cdot D_{z^{k}}V$ (114)
in the constraint of 83, where $B$ and $V$ are the approximately optimal values returned by the SDP solver for the sphere projection method. This expression must vanish on average over the optimal orbit, and it is everywhere non-negative, so close to the orbit it must be very small in value. The minimization strategy is to plot, at each $x_{1}$ and $x_{2}$ over the approximate range of the chaotic attractor, the minimum of this expression over all $z^{k}\in\mathbb{S}^{n_{k}-1}$, which is approximated by random sampling on the unit sphere. This minimum value is then passed to a Newton-shooting algorithm to converge the true periodic orbit, which approximately intersects the point
$(x_{1},x_{2},x_{3},x_{4})=(-0.14984,0.01520,0,1)$ (115)
@@
and has a period of $2\pi$. This is the small, approximately circular periodic orbit depicted in figure 1. We computed LEs on this orbit to be approximately
$\mu_{1}=0.88785,\quad\mu_{2}=0,\quad\mu_{3}=0,\quad\mu_{4}=-1.08785,$ (116)
giving
$M_{1}=M_{2}=M_{3}=0.88785,\quad M_{4}=-0.2.$ (117)
The Kaplan-Yorke formula 18 then gives $d_{L}=3.81615$, given that the supremum must be attained on this orbit. These values perfectly match the bounds given by $C_{1}$, $C_{2}$, $C_{3}$ and $C_{4}$. In the limit $\gamma\to 0$ of 107, the orbit becomes an equilibrium point at the origin. This matches previous observations that global Lyapunov dimension is attained on equilibria in systems that have such points. Finally, we note that our bounds on $d_{L}$ given by $C_{2}$ and $C_{4}$ of 108 are equivalent to $d_{L}\leq 1.8162$ for 107, a significant improvement over the bound of Kuznetsov and Reitmann [18], and that this result must be sharp as an equivalent periodic orbit exists within the nonautonomous system 107.

### v2

@@ 3 Computational examples
We now present computational examples in which our bounding formulations that use SOS optimization are applied to particular ODEs. It has been conjectured that, for all systems with equilibria embedded in chaotic attractors, the supremum in the formula 18 for Lyapunov dimension is attained on an equilibrium [10, 23]. This indeed is the case for many simple systems, including Lorenz’s 1963 model [24]. In order to demonstrate the power of our methods in general, we want examples in which LE sums and Lyapunov dimension are not maximized on equilibria. Therefore we choose one dissipative system and one Hamiltonian system that each have no equilibria.
For all computations reported below, Julia code to replicate the results is available at github.com/jeremypparker/Parker_Goluskin_LEs. We used SumOfSquares.jl [21] to reformulate SOS optimization problems as SDPs, along with SymbolicWedderburn.jl [16] for symmetry exploitation. To solve the resulting SDPs, we primarily used Mosek version 11 [31]. Certain SDPs were also solved using Clarabel.jl [15] and Hypatia.jl [4], which gave very similar results with longer runtimes.
@@
In SOS formulations from the sphere projection approach, the bounds $B$ can be minimized as the objective of the SDPs, and the bounds we report are the numerical optima returned by Mosek. In SOS formulations from the shifted spectrum approach, the value of $B$ must be fixed during each SDP solution because it multiplies coefficients of $V$ in the SOS constraints, and the coefficients to be optimized over can appear only linearly. In such computations, if there exists a $V$ satisfying the SOS constraints, the SDP solver reports ‘feasibile’, and otherwise it reports ‘infeasible’. To minimize the $B$ at which bounds are verified in the shifted spectrum approach, we perform a bisection search in $B$, solving SDPs at various fixed $B$ to find the boundary between large $B$ where the SDPs are feasible and small $B$ where they are infeasible. As $B$ approaches this boundary, numerical conditioning of the SDPs worsens, making it hard to precisely identify the infimum over feasible $B$. Confirming that near-optimal bounds are valid calls for rigorous and/or multiple-precision computations [14, 39], but for simplicity we have used only double precision arithmetic in our examples. Guided by [39], we deem a bound feasible only if Mosek reports a dual objective value of less than $10^{-6}$.
In all cases, the maximum total degree of $V$ in the components of $x$ is restricted to some value $\nu_{x}$. In the sphere projection formulations, we also specify the maximum total degree of $V$ in $z^{k}$ as some value $\nu_{z}$. Then the maximum overall degree of each term is $\max\{\nu_{x},\nu_{z}\}$ when bounding $M_{k}$ using $(x,z^{k})$, or it is $\max\{\nu_{x},2\nu_{z}\}$ when bounding $d_{L}$ using $(x,z^{k},z^{k+1})$. In the shifted spectrum formulations, $V$ is quadratic in $w^{k}$ or $w$ when bounding $M_{k}$ or $d_{L}$, respectively, and in the latter case the maximum total degree of $V$ in $y$ is $\nu_{y}$ and in $(x,y)$ is $\max\{\nu_{x},\nu_{y}\}$. Spaces for polynomials other than $V$ are chosen based on the space for $V$. In particular, maximum degrees of $V$ dictate the maximum degrees of the expressions $S$ that are affine in $V$ and that are constrained to lie in a quadratic module 74. Maximum degrees of any polynomials $\sigma_{i}$ and/or $\rho_{j}$ appearing in the quadratic module representation are chosen such that this representation has the same maximum degrees as $S$. The exact polynomial spaces used in each of our computations for both examples can be seen in the accompanying code.
@@
To confirm the sharpness of our numerical bounds, we also computed various periodic orbits and the LE sums along each of these orbits. We used the Tsit5 solver from the OrdinaryDiffEq.jl package [44] to integrate the ODE systems, find periodic orbits and compute LEs. For an orbit with period $T$, the Jacobian $D\phi^{T}$ of the one-period flow map was constructed by automatic differentiation using the ForwardDiff.jl package [45]. Then, the real parts of the Jacobian’s eigenvalues give the LEs $\mu_{k}$. From these $\mu_{k}$ we calculate the periodic orbit’s LE sums $M_{k}$ and Lyapunov dimension.
@@ 3.1 The Duffing Oscillator
The Duffing oscillator is a second-order non-autonomous ODE governing the amplitude $A(t)$ of oscillations that are damped and sinusoidally forced. Different versions exist, but we choose
$\frac{{\rm d}^{2}A}{{\rm d}t^{2}}+\delta\frac{{\rm d}A}{{\rm d}t}+\beta A+\alpha A^{3}=\gamma\cos{\omega t}$ (102)
with parameter values $(\alpha,\beta,\gamma,\delta,\omega)=(1,-1,0.3,0.2,1)$. Numerical integration gives a trajectory that apparently converges to a chaotic attractor. We numerically estimate the Lyapunov exponents on the chaotic attractor using ChaosTools.jl [6, 5]. This gives $\mu_{1}\approx 0.158$ and $\mu_{2}\approx-0.358$, which imply a local Lyapunov dimension of roughly $1.441$. Kuznetsov and Reitmann [19] derived an analytical upper bound on the Lyapunov dimension $d_{L}$ of the global attractor, which gives $d_{L}\leq 1.9910$ at our parameter values. (In their notation, the present values are $\delta=0.1$, $\varepsilon=0.3$ and $\omega=1$.) Note that we have not defined LEs or Lyapunov dimension for non-autonomous systems, although it is straightforward to do so. Instead, we can reformulate 102 as a first-order system whose right-hand side is autonomous and polynomial, so that the SOS methods of section 2.4 are directly applicable.
To obtain a first-order system from the ODE 102 we let $x_{1}=A$ and $x_{2}=\frac{\mathrm{d}}{\mathrm{d}t}{A}$. To obtain an autonomous polynomial system, we employ a standard trick (e.g., as in Parker et al. [38]) and let $x_{3}=\sin{\omega t}$ and $x_{4}=\cos{\omega t}$. Then the $x(t)$ vector evolves according to
$\frac{\mathrm{d}}{\mathrm{d}t}{x}=f(x),\quad\text{where}\quad f(x)=\begin{bmatrix}x_{2}\\ -\delta x_{2}-\beta x_{1}-\alpha x_{1}^{3}+\gamma x_{4}\\ \omega x_{4}\\ -\omega x_{3}\end{bmatrix}$. (103)
Since $x_{3}$ and $x_{4}$ are sine and cosine of the same argument, dynamics are restricted to the semialgebraic set
$\mathcal{B}=\{x\in\mathbb{R}^{4}~:~h(x)=0\},\quad\text{where}\quad h(x)=x_{3}^{2}+x_{4}^{2}-1.$ (104)
Only odd powers of $x$ appear in 103, so the right-hand side is equivariant under the simple sign-change symmetry $\Lambda=-I$, and the domain $\mathcal{B}$ is $\Lambda$-invariant. This induces a corresponding symmetry in the augmented ODE system for each bounding formulation (cf. proposition 2.21), along with the symmetry that is always present (cf. proposition 2.20). In our code, the symmetries of the relevant augmented system are imposed on $V$, and on any other polynomials appearing in Positivstellansätze, as explained in section 2.5.
For the right-hand side of the ODE 103, the Jacobian matrix is
@@
The additive compound matrices of this Jacobian are
@@
$Df^{[4]}(x)$ $=-\delta,$ (108)
and we recall that $Df^{[1]}=Df$ always. In this system, various relations for the LEs can be deduced a priori. The fourth additive compound—i.e., the trace of the Jacobian—is pointwise constant here, which implies that $M_{4}=-\delta$ on every trajectory. Furthermore, every trajectory has two zero LEs: one associated with constant rotation between the sine and cosine variables ($x_{3}$ and $x_{4}$), and another associated with translation along the flow, which applies to all trajectories since there are no equilibria. We expect all trajectories to have a single positive LE $\mu_{1}>0$, so that the four LEs are $(\mu_{1},0,0,-(\delta+\mu_{1}))$. Note that the two zero LEs are created by the reformulation of the second-order ODE 102 as the fourth-order system 103. Directly applying the definition 14 using the time-dependent dynamics in 8, the original system would have only the two LEs $\mu_{1}$ and $(\delta+\mu_{1})$.
The LE relations described above for the system 103 imply that, on each trajectory, the LE sums are related by $M_{1}=M_{2}=M_{3}$. This further implies that the local Lyapunov dimension is $3+M_{1}/(\delta+M_{1})$, which corresponds to $1+M_{1}/(\delta+M_{1})$ for the original second-order equation 102. Therefore, to obtain sharp upper bounds on all maximum LE sums and on the global Lyapunov dimension, it suffices to find a sharp bound on the maximum leading LE $M_{1}$. Nevertheless, in order to test our methods, we computed bounds using our formulations for $M_{1}$, $M_{2}$, $M_{3}$ and $d_{L}$. We do not report computations for $M_{4}$ since $V=0$ already gives the sharp bound $M_{4}\leq-\delta$, due to $Df^{[4]}=-\delta$ holding pointwise in this system.
To formulate SOS optimization problems, we choose finite-dimensional spaces over which to optimize $V$ and the other polynomials appearing in Positivstellansätze. When using proposition 2.16 to bound $M_{k}$ by the sphere projection approach, for instance, these other polynomials include $\rho$ that multiplies $h$ and $\sigma$ that multiplies $|z^{k}|^{2}-1.$ The spaces of these polynomials are restricted by imposing symmetries of the problem as described in section 2.5. They are also restricted by imposing maximum polynomial degrees of $V$ as described in the fourth paragraph of section 3.
Table 1 reports bounds computed using the sphere projection approach. Bounds on the LE sums $M_{k}$ and Lyapunov dimension $d_{L}$ are computed by solving the SOS optimization problems 80 and 83, respectively, for various polynomial degrees $\nu_{x}$ and $\nu_{z}$.
@@
Table 1: Sphere projection approach for the Duffing oscillator in its autonomous formulation 103. The leading LE sums $M_{k}$ are bounded by solving the SOS problem in proposition 2.16, excluding $M_{4}$ since it is trivial in this system (see text). The Lyapunov dimension $d_{L}$ of the global attractor is bounded by solving the SOS problem in proposition 2.17 with $k=3$; bounds for the non-autonomous formulation 102 can be obtained by subtracting 2. The auxiliary function $V$ is restricted to various maximum degrees $(\nu_{x},\nu_{z})$ in $x$ and $z^{k}$, which induce corresponding degrees in other polynomials, as described in the text. The tabulated bounds are rounded to the precision shown, and the underlines indicate digits that agree (after rounding) with the numerically computed values $M_{1}=M_{2}=M_{3}\approx 0.887897$ and $d_{L}\approx 3.816159$ attained by the periodic orbit shown in figure 1.
$\nu_{x}$ $\nu_{z}$ Upper bounds
$M_{1}$ $M_{2}$ $M_{3}$ $d_{L}$
2 0 27.36247 31.82308 29.78229 3.99116
2 14.82952 15.58938 13.93659 3.97731
4 7.14087 7.86722 7.73049 3.96131
4 0 1.36659 1.36676 1.36662 3.87232
2 0.88790 0.88797 0.88790 3.81616
4 0.88790 0.88790 0.88790 3.81616
6 0 1.18428 1.18427 1.18481 3.85544
2 0.88790 0.88797 0.88790 3.81616
The number of variables in the SOS program 80 for $M_{k}$ with $k=1,2,3$ are $4+C_{4}^{k}=8,10,8$, respectively, and the number of variables in the SOS program 83 for $d_{L}$ is $4+C_{4}^{3}+C_{4}^{4}=9$. All bounds converge quickly as polynomial degrees are raised, and our best bounds on each quantity are sharp to at least 4 digits, as we confirm below by finding a periodic orbit whose LE sums and Lyapunov dimension saturate the bounds to this precision. We will see below that the maximizing orbit is geometrically simple, which typically allows for faster convergence as polynomial degrees are raised, relative to examples with more complicated maximizing orbits [14].
We have also computed bounds using the shifted spectrum approach. In this case, the SOS optimization problems solved to bound $M_{k}$ and $d_{L}$ are propositions 2.18 and 2.19, respectively. All expressions are quadratic in the $w$ variables, and we must choose only the total degree in $x$ for $V$ and for other optimized polynomials. Feasibility or infeasibility of the SOS constraints for the chosen polynomial spaces is determined by solving SDPs with various fixed $B$, and the minimum feasible $B$ is sought by a bisection search. With $\nu_{x}\leq 4$, numerical errors are large, and we did not find clear feasibility at any $B$ value. With $\nu_{x}=6$, however, we obtain the bounds
$M_{1}\leq\underline{0.887898}\quad\text{and}\quad d_{L}\leq\underline{3.8161}7,$ (109)
where underlines indicate digits whose sharpness is confirmed by the periodic orbit described below. The SOS problem 91 and 97, which are feasible when $B$ is fixed to the respective bound values in 109 with $\nu_{x}=6$, do not meet our criterion for feasibility when the final digits of the bounds in 109 are decreased by $1$.
Figure 1: A trajectory on the chaotic attractor (thin line) and the unstable $2\pi$-periodic orbit (thick line) for the Duffing oscillator. The figure is the same for the non-autonomous system 102 or its autonomous reformulation 103.
To seek an orbit that maximizes the leading LE $M_{1}$ among all trajectories, and consequently also maximizes the other $M_{k}$ and the Lyapunov dimension, we follow ideas in [50, 20] for using the $V$ found in SOS computations for bounding time averages. Applied here, their observations imply that that, for any $V(x,z^{1})$ giving a nearly sharp bound $B$ on $M_{1}$ via the SOS program 80, the time average of
$B-\Phi_{1}-f\cdot D_{x}V-\ell_{1}\cdot D_{z^{1}}V$ (110)
is small on any orbit that nearly maximizes $M_{1}$. This quantity is also constrained to be nonnegative, so it must be small pointwise on most parts of such orbits. Therefore we seek points in the $(x,z^{1})$ state space where 110 is small and then seek periodic orbits passing near these points.
To find points where 110 is small, we fix $(x_{3},x_{4})=(0,1)$ and consider a mesh of $(x_{1},x_{2})$ over roughly the extent of the chaotic attractor. At each $(x_{1},x_{2})$ pair, we approximate the minimum of 110 over $z^{1}\in\mathbb{S}^{3}$ by random sampling. For the $(x,z)$ at which we find our smallest value of 110, this $x$ is used as the initial guess for a Newton-shooting algorithm, which converges to a $2\pi$-periodic orbit that approximately intersections the point
$(x_{1},x_{2},x_{3},x_{4})=(-0.14984,0.01520,0,1).$ (111)
@@
This is the small, approximately elliptical periodic orbit in the center of figure 1. We computed LEs on this orbit to be approximately
$\mu_{1}=0.887897,\quad\mu_{2}=0,\quad\mu_{3}=0,\quad\mu_{4}=-1.08785.$ (112)
These LEs give
$M_{1}=M_{2}=M_{3}=0.887897,\quad M_{4}=-0.20000,\quad d_{L}=3.816159,$ (113)
where $d_{L}$ is computed from the Kaplan–Yorke formula 18 with the assumption that $M_{1}$ is indeed maximized on this periodic orbit. The values in 113 match at least 4 digits of our best upper bounds from both the sphere projection approach in table 1 and the shifted spectrum approach in 109. This confirms sharpness of these bounds and maximality of the period orbit in figure 1 to at least 4 digits. In the limit $\gamma\to 0$ of 102, the $2\pi$-periodic orbit becomes an equilibrium at the origin. This is consistent with previous observations that Lyapunov dimension is maximized on equilibria in systems that have such points.
For the Duffing oscillator in its original non-autonomous formulation 102, the sharp bound $d_{L}\leq 1.8162$ on global Lyapunov dimension follows from the bound $d_{L}\leq 3.8162$ that we computed for the autonomous reformulation 103. This improves on the analytical bound from [19], which gives $d_{L}\leq 1.9910$ at the present parameters.